OpenStudy (across):

Suppose you have a water tank and the water in it is being warmed at a constant rate. Additionally, suppose it takes the water a certain amount of time to reach a certain temperature. How would you go about modeling the relationship in between the time it takes the water to reach a certain temperature and its volume? In other words, how is the water volume proportional to the time it takes for it to reach a certain temperature whilst it's being warmed at a constant rate?

OpenStudy (amistre64):

the amount of time it takes to heat up a given amount of water is directly porportional to the volume

OpenStudy (across):

Is it a linear proportion?

OpenStudy (amistre64):

V = h*t perhaps? i dont know about the terminology

OpenStudy (amistre64):

newton has a heat equation i believe :)

OpenStudy (across):

I thought perhaps the heat equation played a role in this case: \[\frac{∂u}{∂t}-\alpha(\frac{∂^2u}{∂x^2}+\frac{∂^2u}{∂y^2}+\frac{∂^2u}{∂z^2})=0\]Maybe I'm just making this way more complicated than it should!

OpenStudy (anonymous):

amistre did u read estudier's questions??

OpenStudy (amistre64):

ive read some of them, but not all of them :)

OpenStudy (anonymous):

did u read "Explain why moderators are an endangered species"

OpenStudy (anonymous):

Lol

OpenStudy (amistre64):

that one? no :)